# Facility Location Games with Scaling Effects

### v2

@@ 7 Approximation Ratio Results
For simplicity of notation, we let $r_{q}:=\frac{\max_{y\in X}q(y)}{\min_{y\in X}q(y)}$ be the ratio between the highest and lowest values of the scaling function. As we will show, our approximation ratio results are typically a function of $r_{q}$. We show that when the scaling function is chosen such that agents have single-peaked preferences, the value of $r_{q}$ is upper bounded by a constant.
@@
When agents’ preferences are single-peaked, $r_{q}\leq e$.
@@
Suppose the scaling function $q$ is a piecewise linear function with $k$ line segment pieces. When agents’ preferences are single-peaked, $r_{q}\leq(1+\frac{1}{k})^{k}$.
@@
When the amount of line segments $k$ approaches infinity, the upper bound of $r_{q}$ asymptotically meets the upper bound of $r_{q}=e$ for general scaling functions. Intuitively, this can also be seen by constructing a piecewise linear function which approximates the function $q(y)=e^{y}$.
@@
We next prove a lemma which, given a scaling function $q$, characterizes phantom mechanisms with scaling that have the best approximation ratio among all such mechanisms which have at least one phantom in $(0,1)$. This statement holds for both objectives of total and maximum cost.
Lemma 2
Let $f$ be a phantom mechanism with scaling, where $p_{1}=0$ and $p_{n+1}=1$. For any given continuous scaling function $q$ with a global minimum in $(0,1)$, if $p_{i}\in(0,1)$ for some $i\in\{2,\dots,n\}$, then there exists a phantom mechanism with scaling $f^{\prime}$ that places the $n-1$ phantoms at a global minimum of the scaling function and has a weakly lower approximation ratio.
Furthermore, for any given continuous scaling function $q$ which only has global minima in $\{0,1\}$, if $p_{i}\in(0,1)$ for some $i\in\{2,\dots,n\}$, then for sufficiently small $\epsilon>0$, there exists a phantom mechanism with scaling $f^{\prime}$ that places the $n-1$ phantoms at $\epsilon$ or $1-\epsilon$ and has a weakly lower approximation ratio.
@@
Throughout the proof, we are primarily concerned with the placement of the $n-1$ phantoms $p_{2},\dots,p_{n}$. We show that for every phantom mechanism with scaling $f$ where at least one phantom $p_{j}$ on $(0,1)$ is not at the lowest point of the scaling function, there exists a different phantom mechanism with scaling $f^{\prime}$ that has a strictly lower approximation ratio. This mechanism places the $n-1$ phantoms at the global minimum of $q$ if it lies in $(0,1)$. Specifically, if the scaling function has multiple global minima, the phantoms are placed at the leftmost global minimum on $(0,1)$. If the scaling function only has global minima on 00 and/or $1$, then for sufficiently small $\epsilon>0$, $f^{\prime}$ places the phantoms at $\epsilon$ or $1-\epsilon$.
For every agent location profile $\boldsymbol{x}$ such that $f(\boldsymbol{x})=p_{j}$, we claim there exists another profile $\boldsymbol{x}^{\prime}$ where, under $f^{\prime}$, the total cost ratio with the optimal total cost is strictly lower. We derive $\boldsymbol{x}^{\prime}$ from $\boldsymbol{x}$ by first moving each agent closer to $f(\boldsymbol{x})=p_{j}$ by a sufficiently small constant factor $c>0$ (so that each agent $i$’s distance changes from $|x_{i}-p_{j}|$ to $c|x_{i}-p_{j}|$), and then translating the profile such that the facility is co-located [^5] This occurs as long as there are agents on both sides of the phantoms. Since the phantoms are placed at $\epsilon$ and $1-\epsilon$ instead of 00 and $1$, there is space on both sides of the phantoms. with the $n-1$ phantoms of $f^{\prime}$ and the relative distances between the agents and the facility is the same but multiplied by the sufficiently small constant factor. The distances are still $c|x_{i}-p_{j}|$, so the ratio of total/maximum unscaled distances is still the same. But now the facility is placed at a point where the scaling function value is strictly lower, so the total/maximum cost ratio is lower. Therefore $f^{\prime}$ has the lowest approximation ratio of all phantom mechanisms with scaling which have at least one phantom on $(0,1)$. ∎
7.1 Total Cost
We now prove approximation ratio results for the objective of total cost, a standard utilitarian measure of efficiency. We begin by giving a lower bound on the approximation ratio achievable by a phantom mechanism, with a proof which applies to any continuous scaling function.
Theorem 7.3
Any phantom mechanism with scaling has a total cost approximation ratio of at least $r_{q}$.
Proof
Given a scaling function $q$, we consider two cases: either all phantoms are located at 00 or $1$, or there is at least one phantom that is not located at 00 or $1$. We start with the former case, where the facility is located at some agent. For every such phantom mechanism with scaling, we can always construct an agent location profile where the scaling function is highest at the agent that receives the facility, so the corresponding total cost must be at least $\max_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, where $f_{med}$ is the median mechanism. Now suppose that the agent location profile has an even number of agents, with $x_{n/2}\neq x_{n/2+1}$, and that the location profile is constructed such that the scaling function is lowest at $(x_{n/2}+x_{n/2+1})/2$. Placing the facility at this point results in the optimal total cost of $\min_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, and thus the total cost approximation ratio must be at least $r_{q}$ in this case.
Now consider the case where at least one phantom is located in $(0,1)$. From Lemma 2, it suffices to consider mechanism $f^{\prime}$, a phantom mechanism with scaling which places one phantom on 00, one phantom on $1$, and the remaining $n-1$ phantoms on the leftmost global minimum of $q$ if there exists a global minimum of $q$ in $(0,1)$. If all global minima of $q$ are in $\{0,1\}$, then for sufficiently small $\epsilon>0$, the $n-1$ phantoms are placed on $\epsilon$ or $1-\epsilon$.
We show that $f^{\prime}$ has an approximation ratio of at least $r_{q}$. We define a scaling function $q$ which has a unique global minimum at some point $c_{1}\in(0,1)$, a unique global maximum at $c_{2}>c_{1}$, and a point $c_{3}>c_{2}$ where $q(c_{3})=q(c_{1})+\epsilon$, for sufficiently small $\epsilon>0$. Consider a location profile where $x_{1}=c_{2}<x_{2}=x_{3}=\dots=x_{n}=c_{3}$. The mechanism $f^{\prime}$ places the facility at $x_{1}$, resulting in a total cost of $(n-1)(x_{n}-x_{1})\max_{y}q(y)$, whilst the optimal total cost is at most $(x_{n}-x_{1})(\min_{y}q(y)+\epsilon)$ by placing the facility at $c_{3}$. Therefore $f^{\prime}$ has an approximation ratio of at least $r_{q}$. By exhaustion of cases, we see that any phantom mechanism has a total cost approximation ratio of at least $r_{q}$. ∎
@@ 7.1 Total Cost
From this result, we can also easily see that among all of the phantom mechanisms, placing the facility at the median agent gives the best possible approximation ratio for total cost.
@@ 7.2 Maximum Cost
We next prove approximation ratio results for the objective of maximum cost, a standard measure of egalitarian fairness. We begin by giving the maximum cost approximation ratio of any phantom mechanism with scaling with $p_{1}=0$ and $p_{n+1}=1$, and again the proof applies to any arbitrary continuous scaling function.
@@
Any phantom mechanism with scaling has a maximum cost approximation ratio of $2r_{q}$.
@@
Similar to the proof of Theorem 7.3, given a scaling function $q$, we consider two cases: either all phantoms are located at 00 or $1$, or there is at least one phantom that is not located at 00 or $1$. In the former case, where the facility is located at some agent, we can always construct an agent location profile where the scaling function is highest at the agent that receives the facility, so the corresponding maximum cost must be at least $\max_{y}q(y)\sum_{i}|f_{mid}(\boldsymbol{x})-x_{i}|$, where $f_{mid}$ is the midpoint mechanism. Now suppose that the location profile is also constructed such that the scaling function is lowest at $(x_{1}+x_{n})/2$. Placing the facility at this point results in the optimal maximum cost of $\min_{y}q(y)\sum_{i}|f_{mid}(\boldsymbol{x})-x_{i}|$, and thus the total cost approximation ratio must be at least $2r_{q}$ in this case.
For the case where at least one phantom is not located at 00 or $1$, by Lemma 2, it suffices to consider the phantom mechanism with scaling $f^{\prime}$, which is defined in Lemma 2 and the proof of Theorem 7.3. We show that the approximation ratio of $f^{\prime}$ is at least $2r_{q}$. Consider a scaling function with two global minima at $c_{1}$ and $c_{3}$, and a global maximum at $c_{2}$, where $0<c_{1}<c_{2}<c_{3}$. The $n-1$ phantoms of $f^{\prime}$ are placed at $c_{1}$. We construct an agent location profile such that $x_{1}=\dots=x_{n-1}=c_{2}$, and $x_{n}=2c_{3}-c_{2}$. Here, $f^{\prime}$ places the facility at $c_{2}$, giving a maximum cost of $\max_{y\in X}q(y)(2c_{3}-2c_{2})$. However, the optimal facility placement is at $c_{3}$, which gives a maximum cost of $\min_{y\in X}q(y)(c_{3}-c_{2})$. Dividing these terms gives us the lower bound of $2r_{q}$.
To see that the approximation ratio is at most $2r_{q}$, note that a phantom mechanism with $p_{1}=0$ and $p_{n+1}=1$ ensures the facility is placed within $[x_{1},x_{n}]$, so the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$, whilst the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$. ∎

### v3

@@ 7 Approximation Ratio Results
For simplicity of notation, we let $r_{q}:=\frac{\max_{y\in X}q(y)}{\min_{y\in X}q(y)}$ be the ratio between the highest and lowest values of the scaling function. As we will show, our approximation ratio results are typically a function of $r_{q}$. We show that when the scaling function is chosen such that agents are guaranteed to have single-peaked preferences, the value of $r_{q}$ is upper bounded by a constant.
@@
When agents are guaranteed to have single-peaked preferences, $r_{q}\leq e$.
@@
Suppose the scaling function $q$ is a piecewise linear function with $k$ line segment pieces. When agents are guaranteed to have single-peaked preferences, $r_{q}\leq(1+\frac{1}{k})^{k}$.
@@
When the number of line segments $k$ approaches infinity, the upper bound of $r_{q}$ asymptotically meets the upper bound of $r_{q}=e$ for general scaling functions. Intuitively, this can also be seen by constructing a piecewise linear function which approximates the function $q(y)=e^{y}$.
@@
7.1 Total Cost
We now prove approximation ratio results for the objective of total cost, a standard utilitarian measure of efficiency. We begin by giving a lower bound on the approximation ratio achievable by a phantom mechanism with scaling.
Theorem 7.3
There exists a scaling function $q$ such that any phantom mechanism with scaling has a total cost approximation ratio of at least $r_{q}$.
@@ 7.1 Total Cost
By Lemma 1, it suffices to consider phantom mechanisms with scaling where at least one phantom is on 00 and at least one phantom is on $1$ (regardless of the scaling function), so for the remainder of the proof, we are concerned with the placement of the remaining $n-1$ phantoms.
Consider an arbitrary phantom mechanism with scaling $f$, and a continuous scaling function $q$ which has global maxima at 00, $0.5$, and $1$, and global minima at $0.25$ and $0.75$. We divide the remainder of the proof into two cases. In the first case, suppose that under $q$, $f$ places $k\in\{0,\dots,n-2\}$ phantoms in $[0,0.5]$ (and the remaining phantoms in $(0.5,1]$). Then for an agent location profile with even $n$, where $\frac{n}{2}$ agents are on 00 and $\frac{n}{2}$ agents are on $0.5$, the facility will be placed on $0.5$. This results in a total cost of $\max_{y}q(y)\frac{n}{2}(0.5-0)$, but the optimal facility placement of $0.25$ gives a total cost of $\min_{y}q(y)*0.25n$. Dividing these terms gives us our lower bound of $r_{q}$. In the remaining case, suppose that under $q$, $f$ places all $n-1$ phantoms in $[0,0.5]$. Then we consider an agent location profile with even $n$ where $\frac{n}{2}$ agents are on $0.5$ and $\frac{n}{2}$ agents are on $1$. Here, the facility will be placed on $0.5$ to give a total cost of $\max_{y}q(y)\frac{n}{2}(1-0.5)$, but the optimal facility placement of $0.75$ gives a total cost of $\min_{y}q(y)*0.25n$. Again, dividing these terms gives our lower bound of $r_{q}$, completing the proof by exhaustion of cases. ∎
@@
From this result, we can also easily see that among all of the phantom mechanisms with scaling, simply placing the facility at the median agent gives the best possible approximation ratio for total cost.
@@ 7.2 Maximum Cost
We next prove approximation ratio results for the objective of maximum cost, a standard measure of egalitarian fairness. We begin by giving a lower bound for the maximum cost approximation ratio of any phantom mechanism with scaling.
@@
There exists a scaling function $q$ such that any phantom mechanism with scaling has a maximum cost approximation ratio of at least $2r_{q}$.
@@
The proof of this theorem is almost identical to the proof of Theorem 7.3; we repeat the key arguments for completeness. By Lemma 1, it suffices to consider mechanisms which always place at least one phantom on 00 and at least one phantom on $1$, so we are concerned with the placement of the remaining $n-1$ phantoms.
For the instance described in the proof of Theorem 7.3, if at most $n-2$ phantoms are placed in $[0,0.5]$, the maximum cost from the facility placement at $0.5$ will be $\max_{y}q(y)(0.5-0)$, whilst the optimal maximum cost is $\min_{y}q(y)(0.5-0.25)$. By dividing these terms, we obtain our lower bound of $2r_{q}$. In the remaining case where all $n-1$ phantoms are placed in $[0,0.5]$, the maximum cost from the facility placement is $\max_{y}q(y)(1-0.5)$, whilst the optimal maximum cost is $\min_{y}q(y)(0.75-0.5)$. Again, dividing these terms gives our lower bound of $2r_{q}$, completing the proof. ∎
Unlike the total cost objective, we find that under the maximum cost objective, any phantom mechanism with scaling which always places at least one phantom on 00 and at least one phantom on $1$ has the same maximum cost approximation ratio which matches our lower bound.
Theorem 7.6
Any phantom mechanism with scaling which, for any $q\in\mathcal{Q}$, places at least one phantom on 00 and at least one phantom on $1$, has a maximum cost approximation ratio of $2r_{q}$.
Proof
By Theorem 7.5, it suffices to show that the approximation ratio is at most $2r_{q}$. To see this, note that under a phantom mechanism with scaling which places at least one phantom on 00 and at least one phantom on $1$, the facility is always placed within $[x_{1},x_{n}]$. As a result, the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$, whilst the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$. ∎